The effectiveness of a blood-pressure drug is being investigated. An experimenter finds that, on average, the reduction in systolic blood pressure is 70.2 for a sample of size 25 and standard deviation 17.7. Estimate how much the drug will lower a typical patient's systolic blood pressure (using a 98% confidence level). Assume the data is from a normally distributed population.

Enter your answer as a tri-linear inequality accurate to three decimal places.

__________< μ < ______
Answer #1

Solution :

Given that,

sample size = n = 25

Degrees of freedom = df = n - 1 = 25 - 1 = 24

t /2,df = 2.492

Margin of error = E = t/2,df * (s /n)

= 2.492 * (17.7 / 25)

Margin of error = E = 8.822

The 98% confidence interval estimate of the population mean is,

- E < < + E

70.2 - 8.822 < < 70.2 + 8.822

61.378 < < 79.022